Happy Friday to you math pros! Can I get a little help from you this morning?
The expression for figuring a student's average test score is p/t, where p is the total points earned on all the tests, and t is the number of tests. Find a student's average when 344 points are earned on 4 tests.
The average is____ points.
thanks

Answer by stanbon(75887)

344/4 = 172/2 = 86
====================
Cheers,
Stan H.